# Bilateral Distribution Compression: Reducing Both Data Size and Dimensionality

### v3

@@ 2. Step 2:
    Given the optimal encoder $\psi_{\bm{\alpha}_{*}}$, optimise the compressed set $\mathcal{C}\subset\mathbb{R}^{p}$ to minimise the EJMMD. That is, we solve
@@
Corollary C.1
Given $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, $h:\mathbb{R}^{p}\times\mathbb{R}^{p}\to\mathbb{R}$, and $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$, ket $r^{d}\left((\bm{x},\bm{y}),(\bm{x}^{\prime},\bm{y}^{\prime})\right)=k(\bm{x},\bm{x}^{\prime})\,l(\bm{y},\bm{y}^{\prime})$ and $r^{p}\left((\bm{z},\bm{y}),(\bm{z}^{\prime},\bm{y}^{\prime})\right)=h(\bm{z},\bm{z}^{\prime})\,l(\bm{y},\bm{y}^{\prime})$ be characteristic tensor product kernels. Further, assume $\phi\circ\psi:\mathcal{X}\to\mathcal{X}$ is an autoencoder defined by the composition of measurable maps $\psi:\mathcal{X}\to\mathbb{R}^{p}$ and $\phi:\mathbb{R}^{p}\to\mathcal{X}$. Suppose $\mathbb{P}_{X,Y}$ and $\mathbb{P}_{Z,W}$ are distributions such that
@@
C.4 Complexity Analysis
@@
which is linear in the dataset size $n$ and the input dimension $d$. Note that by ignoring constants $E,B,T$ and noting $p\ll d$ and $m\ll n$, we have that BDC is overall $\mathcal{O}(nd)$.
@@ Appendix D Experiment Details
For all experiments, we use the Adam optimiser (Kingma & Ba, 2017) as a sensible default choice. However, the implementations allow for an arbitrary choice of optimiser via the Optax package (DeepMind et al., 2020).
@@ D.1.2 Additional Regression Experiment: Swiss-Roll with Linear Projection and IMQ Kernel
In this section, we revisit the linear Swiss-Roll dataset from Section D.1.1, this time using the inverse multi-quadratic (IMQ) kernel instead of the Gaussian kernel for distribution compression. IMQ kernels are characteristic (Weinstein & Miller, 2021) and take the form:
@@ D.2 Exact Bilateral Distribution Compression
$\mu_{\mathbb{P}_{X}}(\bm{x})$ $=\left|I+\tfrac{1}{\lambda^{2}}\Sigma\right|^{-1/2}\exp\left(-\tfrac{1}{2}(\bm{x}-\bm{\mu})^{\top}\big(\lambda^{2}I+\Sigma\big)^{-1}(\bm{x}-\bm{\mu})\right),$
$\mathbb{E}_{\mathbb{P}_{X}}\left[\mu_{\mathbb{P}_{X}}(X)\right]$ $=\left|I+\frac{2}{\lambda^{2}}\Sigma\right|^{-1/2}$
@@
$\mu_{\mathbb{P}_{X}}(\bm{x})$ $=\sum_{i}\omega_{i}\left|I+\tfrac{1}{\lambda^{2}}\Sigma_{i}\right|^{-1/2}\exp\left(-\tfrac{1}{2}(\bm{x}-\bm{\mu}_{i})^{\top}\big(\lambda^{2}I+\Sigma_{i}\big)^{-1}(\bm{x}-\bm{\mu}_{i})\right),$
$\mathbb{E}_{\mathbb{P}_{X}}\left[\mu_{\mathbb{P}_{X}}(X)\right]$ $=\sum_{i,j}\omega_{i}\omega_{j}\left|I+\frac{\Sigma_{i}+\Sigma_{j}}{\lambda^{2}}\right|^{-\frac{1}{2}}\exp\left(-\tfrac{1}{2}(\bm{\mu}_{i}-\bm{\mu}_{j})^{\top}\big(\lambda^{2}I+\Sigma_{i}+\Sigma_{j}\big)^{-1}(\bm{\mu}_{i}-\bm{\mu}_{j})\right).$
@@ D.3 Decoded MNIST Images
In Figure 20, the reconstructed images produced by BDC-NL are sharper than those from BDC-L, which is expected since the nonlinear autoencoder can recover more information than the linear variant. Comparing M3D and ADC, the M3D images also appear sharper and the sharpness of the M3D reconstructions is broadly comparable to that of BDC-NL. While these images are useful for visual comparison and help us reason about the different methods, it is important to note that the downstream task in BDC is performed directly in the latent space rather than in the ambient image space. Indeed, in Figure 7 we observed that BDC-L achieved the best Gaussian Process predictive metrics in the shortest time, despite producing lower-quality reconstructions here. This suggests that the latent space learned by BDC-L may be more informative about the feature–response relationship than that of BDC-NL. Moreover, both BDC methods substantially outperformed ADC and M3D.

### v4

@@ 2. Step 2:
    Given the optimal encoder $\psi_{\bm{\alpha}_{*}}$, optimise the compressed set $\mathcal{C}\subset\mathbb{R}^{p}\times\mathcal{Y}$ to minimise the EJMMD. That is, we solve
@@
Corollary C.1.
Given $k:\mathcal{X}\times\mathcal{X}\to\mathbb{R}$, $h:\mathbb{R}^{p}\times\mathbb{R}^{p}\to\mathbb{R}$, and $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$, let $r^{d}\left((\bm{x},\bm{y}),(\bm{x}^{\prime},\bm{y}^{\prime})\right)=k(\bm{x},\bm{x}^{\prime})\,l(\bm{y},\bm{y}^{\prime})$ and $r^{p}\left((\bm{z},\bm{y}),(\bm{z}^{\prime},\bm{y}^{\prime})\right)=h(\bm{z},\bm{z}^{\prime})\,l(\bm{y},\bm{y}^{\prime})$ be characteristic tensor product kernels. Further, assume $\phi\circ\psi:\mathcal{X}\to\mathcal{X}$ is an autoencoder defined by the composition of measurable maps $\psi:\mathcal{X}\to\mathbb{R}^{p}$ and $\phi:\mathbb{R}^{p}\to\mathcal{X}$. Suppose $\mathbb{P}_{X,Y}$ and $\mathbb{P}_{Z,W}$ are distributions such that
@@
C.4 Practitioners Guide to Choosing Hyperparameters
The algorithms developed in this work involve several hyperparameters that must be selected and tuned, namely the choice of kernel functions $k(\cdot,\cdot)$ and $h(\cdot,\cdot)$, the compressed set size $m$, and the latent dimension $p$.
The choice of kernel primarily reflects the practitioner’s inductive bias, and there is extensive literature on kernel selection; see, for example, Hofmann et al. (2008). In practice, however, many commonly used kernels yield comparable performance when reasonably parametrised. In Section D.1.2, we ablate over kernel choices on the Swiss-Roll dataset and observe no degradation in the relative performance of the method (Figure 14), indicating robustness to this choice in our setting.
The selection of the compressed set size $m$ is driven largely by computational considerations. Larger values of $m$ improve fidelity to the target distribution but incur higher computational cost. In Section D.1.8, we study downstream task performance as a function of $m$, with Figure 23 showing the expected trade-off: increasing $m$ improves performance while increasing runtime requirements. In practice, $m$ should therefore be chosen to balance computational constraints against accuracy requirements.
The latent dimension $p$ plays a different role, as it determines the geometry in which downstream tasks operate. If $p$ is too small, important structure may be lost through excessive compression, whereas overly large values of $p$ can lead to unnecessary optimisation difficulty without corresponding gains. Fortunately, there is extensive work on intrinsic dimension estimation, and practical estimators are readily available (Bac et al., 2021). In Section D.1.6, we perform an ablation study on the choice of $p$ on the CT-SLice dataset, finding that the optimal latent dimension for both BDC-L and BDC-NL is approximately $p\approx 10$. Applying the suite of intrinsic-dimension estimators from the skdim package (Bac et al., 2021) yields an average estimate of approximately $11.5$, indicating close agreement between empirically optimal performance and standard dimensionality estimates. Throughout this work, we primarily selected $p$ based on computational limitations and desired compression rates. However, the results of Section D.1.6 demonstrate that this strategy is not necessarily optimal when downstream performance is the primary objective, and that independent tuning of $p$ via skdim can yield further improvements.
C.5 Complexity Analysis
@@
which is linear in the dataset size $n$ and the input dimension $d$.
In practice, the number of training epochs $E$ is small; across all experiments we used at most $E=100$. The batch size $B$ is treated as a constant and never exceeds $B=1024$ here, and optimisation in the latent space converges rapidly, with no experiment requiring more than T=5, [^0] T=5{,}000 gradient steps. Since $p\ll d$ and $m,E,B,T\ll n$, the total computational cost of BDC scales linearly in the ambient data size. In particular, for fixed $m$, $p$, $B$, $E$, and $T$, the overall complexity is $\mathcal{O}(nd)$.
@@ Appendix D Experiment Details
For all experiments, we use the Adam optimiser (Kingma and Ba, 2017) as a sensible default choice. However, the implementations allow for an arbitrary choice of optimiser via the Optax package (DeepMind et al., 2020).
@@ D.1.2 Additional Regression Experiment: Swiss-Roll with Linear Projection and IMQ Kernel
In this section, we revisit the linear Swiss-Roll dataset from Section D.1.1, this time using the inverse multi-quadratic (IMQ) kernel instead of the Gaussian kernel for distribution compression. IMQ kernels are characteristic (Weinstein and Miller, 2021) and take the form:
@@ D.1.7 Effect of Increasing Size of Latent Dimension on MNIST
D.1.8 Effect of Increasing Size of the Compressed Set on CT-Slice
In this section, we investigate how increasing the number of observations in the compressed set affects GP regression performance on the CT-Slice dataset from Section 4.1.1. We use the same experimental setup, however we now construct compressed sets up to size $m=500$.
Figure 23 shows that increasing the size of the compressed set consistently improves performance, as expected. However, we also observe diminishing returns, particularly for the mean squared error. For the uncertainty-focused metrics (NLL and CRPS), the drop-off is less pronounced, which suggests that larger compressed sets play a more important role in achieving accurate uncertainty quantification.
Across all compressed-set sizes, the bilateral methods deliver a substantial performance gain over ADC. Although constructing larger compressed sets naturally increases computation time, ADC remains significantly more expensive than the bilateral methods, and this gap widens as the compressed-set size grows. This is intuitive: for the bilateral methods, the dominant cost lies in learning the latent space. Once that space is learned, gradient descent in a low-dimensional setting is very fast.
Figure 23: GP performance versus size of compressed set on the CT-Slice dataset. We report ADC (red), BDC with a nonlinear autoencoder (BDC-NL, green) and BDC with a linear autoencoder (BDC-L, red), averaged over three runs, as well as URS (grey, shaded bands: 25th-75th percentiles) over 100 runs.
D.1.9 Effect of Weighting the Hybrid Loss
In Section 3.2.2 we introduced an MSRE term to the reconstruction loss for nonlinear autoencoders in Equation 5. This was included to avoid a failure mode characterised by catastrophic mixing. The RMMD term does not guard against this behaviour for highly flexible autoencoders. In the main body we weighted these terms equally. All datasets are normalised prior to training and the MSRE term is further normalised by $\frac{1}{nd}$. This guarantees that the MSRE component of the loss does not grow arbitrarily with the dataset size or dimensionality and it places RMMD and MSRE on comparable scales. For this reason we chose an equal weighting between the two terms. This performed well empirically and avoided the need to introduce an additional hyperparameter that could complicate optimisation.
However, it is natural to introduce a weighting parameter, and to investigate any potential trade-offs. We therefore replace Equation 5 with
$\mathrm{RMMD}_{k}^{2}\left(\hat{\mathbb{P}}_{X},\hat{\mathbb{P}}_{\phi(\psi(X))}\right)+\frac{\lambda}{nd}\sum_{i=1}^{n}\lVert\mathbf{x}_{i}-\phi(\psi(\mathbf{x}_{i}))\rVert_{2}^{2},$ (8)
where $\lambda>0$ is a real-valued weight that determines the relative strength of the MSRE term with respect to the RMMD term.
In Figure 24 we report downstream GP task performance on the CT-Slice dataset for values of $\lambda$ ranging from $0.1$ to $10$, which represents a change of two orders of magnitude in the importance of the MSRE term. Once the RMMD term is sufficiently weighted to prevent mixing collapse, downstream performance remains relatively stable across this broad range of MSRE weights. This suggests that the method is not highly sensitive to the relative weighting and that the equal-weight choice is adequate for this case. Interestingly we also see that the equal weighting
Figure 24: GP test performance and test RMMD as functions of the MSRE weighting strength on the CT-Slice dataset. We report results for BDC with a nonlinear autoencoder (BDC-NL, green), with shaded bands indicating the 25th–75th percentile range. The vertical black dashed line denotes the point of equal weighting.
The weighting formulation in Equation 8 suggests a potentially promising direction for BDC optimisation with neural autoencoders. Adaptive loss balancing methods such as those introduced in Chen et al. (2018) treat the weight $\lambda$ as a learnable parameter that is updated during optimisation. It is possible that such a method could lead to improved results for BDC-NL.
D.1.10 Comparison to a Sequential Baseline
A natural baseline for bilateral distribution compression is a sequential pipeline that first performs MMD-based compression in the ambient space (via ADC), and subsequently applies dimensionality reduction such as PCA or an autoencoder. While conceptually straightforward, this approach is both computationally inefficient and statistically suboptimal compared to the integrated formulation proposed in BDC.
From a computational perspective, performing MMD-based optimisation in the ambient space is expensive. As shown in Table 2, gradient-based compression in high dimensions incurs a substantial cost. BDC avoids this bottleneck by learning a low-dimensional latent representation prior to compression, after which MMD-based coreset construction is performed in the latent space.
More importantly, a sequential approach fundamentally constrains the quality of the learned latent representation. Any information discarded during the initial ambient-space compression step is irrevocably lost, limiting the expressiveness of the subsequent dimensionality reduction. Since the geometry of the latent space is critical for downstream tasks, it is essential that it be learned directly from the full dataset rather than from a compressed proxy.
To make this failure mode explicit, we consider an additional baseline in which ADC is first applied in the ambient space, followed by PCA on the resulting compressed set. Figure 25 reports results on the CT-Slice dataset. Both variants of BDC significantly outperform this sequential baseline across compressed set sizes, while also being considerably faster to compute.
Figure 25: GP performance versus compressed set size on the CT-Slice dataset. We report ADC (purple), BDC with a nonlinear autoencoder (BDC-NL, green), and BDC with a linear autoencoder (BDC-L, red), averaged over three runs. Uniform random sampling (URS, grey) is shown with shaded bands indicating the 25th to 75th percentiles over 100 runs.
Finally, standard dimensionality reduction techniques such as PCA do not provide distributional guarantees. As shown in Theorem 3.1, PCA preserves only second-order structure. In contrast, BDC leverages characteristic kernels and directly optimises distributional objectives, ensuring preservation of all moments of the target distribution under Theorems 3.3 and 3.5. This distinction is crucial for probabilistic downstream tasks, where higher-order structure directly impacts predictive performance.
@@ D.2 Exact Bilateral Distribution Compression
$\mu_{{X}}(\bm{x})$ $=\left|I+\tfrac{1}{\lambda^{2}}\Sigma\right|^{-1/2}\exp\left(-\tfrac{1}{2}(\bm{x}-\bm{\mu})^{\top}\big(\lambda^{2}I+\Sigma\big)^{-1}(\bm{x}-\bm{\mu})\right),$
$\mathbb{E}_{\mathbb{P}_{X}}\left[\mu_{X}(X)\right]$ $=\left|I+\frac{2}{\lambda^{2}}\Sigma\right|^{-1/2}$
@@
$\mu_{{X}}(\bm{x})$ $=\sum_{i}\omega_{i}\left|I+\tfrac{1}{\lambda^{2}}\Sigma_{i}\right|^{-1/2}\exp\left(-\tfrac{1}{2}(\bm{x}-\bm{\mu}_{i})^{\top}\big(\lambda^{2}I+\Sigma_{i}\big)^{-1}(\bm{x}-\bm{\mu}_{i})\right),$
$\mathbb{E}_{{X}}\left[\mu_{X}(X)\right]$ $=\sum_{i,j}\omega_{i}\omega_{j}\left|I+\frac{\Sigma_{i}+\Sigma_{j}}{\lambda^{2}}\right|^{-\frac{1}{2}}\exp\left(-\tfrac{1}{2}(\bm{\mu}_{i}-\bm{\mu}_{j})^{\top}\big(\lambda^{2}I+\Sigma_{i}+\Sigma_{j}\big)^{-1}(\bm{\mu}_{i}-\bm{\mu}_{j})\right).$
@@ D.3 Decoded MNIST Images
In Figure 26, the reconstructed images produced by BDC-NL are sharper than those from BDC-L, which is expected since the nonlinear autoencoder can recover more information than the linear variant. Comparing M3D and ADC, the M3D images also appear sharper and the sharpness of the M3D reconstructions is broadly comparable to that of BDC-NL. While these images are useful for visual comparison and help us reason about the different methods, it is important to note that the downstream task in BDC is performed directly in the latent space rather than in the ambient image space. Indeed, in Figures 7 and 8 we observed that BDC-L achieved the best Gaussian Process predictive metrics in the shortest time, despite producing lower-quality reconstructions here. This suggests that the latent space learned by BDC-L may be more informative about the feature–response relationship than that of BDC-NL. Moreover, both BDC methods substantially outperformed ADC and M3D.